OpenStudy (anonymous):

Which step should be completed first to solve the equation ? A. Add 12 to both sides. B. Add 13 to both sides. C. Divide both sides by 8. D. Subtract 13 from both sides.

OpenStudy (chycora):

were is the equation

OpenStudy (anonymous):

oh the equation is 8x-12=13

OpenStudy (anonymous):

so whats the answer

OpenStudy (chycora):

the answer is to add 12 to the 13 like this 12+13

OpenStudy (anonymous):

ok

OpenStudy (chycora):

so do you kn which one it is

OpenStudy (anonymous):

25

OpenStudy (anonymous):

12

OpenStudy (chycora):

yes

OpenStudy (anonymous):

add 12 to both sides

OpenStudy (chycora):

so it is a if i'm correct

OpenStudy (chycora):

do you agree

OpenStudy (anonymous):

thanks:))

OpenStudy (chycora):

welcome

OpenStudy (chycora):

need help with anything else

OpenStudy (chycora):

r u ok

OpenStudy (anonymous):

ya sure wait a minate

OpenStudy (chycora):

ok

OpenStudy (chycora):

wat grade r u in

OpenStudy (anonymous):

What is the solution to the equation ? 8-2b=-2/3(12b+6)? b =

OpenStudy (anonymous):

7th

OpenStudy (chycora):

kool me 2

OpenStudy (anonymous):

help

OpenStudy (anonymous):

please

OpenStudy (anonymous):

r u there

OpenStudy (chycora):

i can help you hold up ok

OpenStudy (anonymous):

k

OpenStudy (anonymous):

chycora

OpenStudy (chycora):

yes

OpenStudy (anonymous):

help?

OpenStudy (chycora):

ok the answer is -1.25 i think and if it's not i am so sorry

OpenStudy (anonymous):

it cant fit in the box that means its rong

OpenStudy (anonymous):

thank u anyways

OpenStudy (chycora):

ok let me think about it

OpenStudy (chycora):

nvm i'm so srry idk srry